# Quantization Error Propagation: Revisiting Layer-Wise Post-Training Quantization

### v2

@@ 5.3 Controlling Propagation Strength
Although solving Eq. (5) effectively reduces the accumulation of quantization error, it can lead to overfitting. This issue is particularly pronounced when the calibration dataset is small, insufficiently representative of the target task, or when the model includes blocks with a large number of parameters such as the MLP blocks commonly found in transformer architectures, causing the correction to overfit the calibration dataset.
@@
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times m}}\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}^{2}+\lambda_{l}\|{\bm{W}}_{l}-\widehat{{\bm{W}}}_{l}\|_{F}^{2},\quad\lambda_{l}\in{\mathbb{R}}_{+}.$ (9)
Specifically, as $\alpha_{l}$ increases from 00 to $1$, the corresponding parameter $\lambda_{l}$ decreases from +$\infty$+ ∞ to 00.
A detailed derivation is provided in Appendix A.4. Additionally, the following proposition is established.
@@
Explicit conditions and comprehensive proofs of this proposition are provided in Appendix A.3. Consequently, the parameter $\alpha_{l}$ effectively controls overfitting, analogous to regularization techniques, and importantly provides a systematic way to balance overfitting and underfitting in layer-wise PTQ methods. Indeed, this parameter is crucial for preventing overfitting, especially in MLP blocks, which contain more parameters compared to other blocks.
Furthermore, in large-scale LLMs, the high dimensional activations in MLP layers often results in computationally expensive correction terms. In these cases, selectively setting $\alpha_{l}=0$ for specific layers eliminates the computational cost of the correction term and acts as implicit regularization, potentially improving generalization. Therefore, appropriately setting $\alpha_{l}=0$ can reduce the correction time by approximately one-third and one-half. Developing adaptive strategies for layer-wise, data-aware, or resource-efficient tuning of $\alpha_{l}$ is a promising direction for future research. In the following, we refer to the overall approach, including the tunable mechanism controlled by $\{\alpha_{l}\}_{l=1}^{L}$, as Quantization Error Propagation (QEP).
@@ 6 Experiments
We use representative layer-wise PTQ methods based on linear quantization such as round-to-nearest (RTN) (Frantar et al., 2022; Dettmers and Zettlemoyer, 2023), GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2024) and QuIP (Chee et al., 2023). Although previous studies have explored extensions, such as non-linear quantization and block-wise quantization, as discussed in Section 2, these techniques are orthogonal to the core improvement introduced by QEP. Therefore, to clearly isolate and emphasize the impact of QEP, we focus on these representative layer-wise PTQ methods.
@@
This study focuses on weight-only quantization schemes, specifically per-channel and group-wise quantization, which have recently shown superior trade-offs between efficiency and accuracy (Dettmers and Zettlemoyer, 2023; Frantar et al., 2022; Lin et al., 2024). In the main text, we evaluates per-channel quantization under INT4, INT3, and INT2 precision settings. Due to space constraints, detailed results for group-wise quantization are presented in Appendix C. For the propagation strength parameter $\alpha_{l}$, we adopt a representative default value of $\alpha_{l}=\nicefrac{{1}}{{2}}$ for all layers, except for the MLP layers in the Llama-2 70B model, for which we set $\alpha_{l}=0$. Tuning $\alpha_{l}$ can further improve performance but is beyond the scope of this study and is left for future work.
@@
Following previous studies, we evaluate the Hessian matrix using the same default calibration datasets as used in their original implementations. Specifically, GPTQ and QuIP use the C4 dataset (Frantar et al., 2022) for calibration, while AWQ uses the Pile dataset (Gao et al., 2020). Following Frantar et al. (2022), we evaluate the correction term in Eq. (4) using $128$ randomly sampled segments of $2048$ tokens each from the C4 dataset(Raffel et al., 2020), which consists of web-crawled text excerpts.
@@
Following Lin et al. (2024); Frantar et al. (2022), we evaluate our method on recent popular LLMs, namely the Llama-2 and Llama-3 model families (Touvron et al., 2023), with size ranging from 7B to 70B parameters, as well as Mistral-7B (Jiang, 2024). These models demonstrate superior performance compared to other open-source LLMs (Zhang et al., 2022; Workshop et al., 2022) and have become widely adopted as foundational models for numerous derivative open-source models (Taori et al., 2023; Chiang et al., 2023).
@@
Following established evaluation protocols from prior studies (Dettmers et al., 2022; Xiao et al., 2023; Frantar et al., 2022; Dettmers and Zettlemoyer, 2023; Yao et al., 2022), we evaluate the quantized LLMs using the perplexity (PPL) on WikiText2 (Merity et al., 2016), Penn Treebank (PTB) (Marcus et al., 1994), and C4 (Raffel et al., 2020), and zero-shot accuracy on benchmarks including ARC Easy (ArcE) (Boratko et al., 2018), PiQA (Bisk et al., 2020), and StoryCloze (SC) (Mostafazadeh et al., 2016). Due to space limitations, detailed results for each dataset are provided in Appendix C. All experiments are conducted using a single NVIDIA V100 GPU.
@@ 6.1 Results
Table 1 summarizes PPL results of various quantized models evaluated on WikiText2, comparing several bit-widths and different layer-wise PTQ methods, both with and without QEP. Additional results on C4 and PTB datasets are provided in Appendix C.1, demonstrating consistent trends in the following. Our results indicate that incorporating QEP significantly enhances the performance of layer-wise PTQ, substantially reducing perplexity across nearly all tested methods and quantization levels. In medium-bit scenarios such as INT4 and INT3, where AWQ already exhibits strong performance, applying QEP yields further improvements. At 2-bit quantization, existing layer-wise PTQ methods based on linear quantization typically suffer sever PPL degradation, rendering deployment infeasible. However, QEP effectively mitigates this issue, making INT2 quantization achievable with practical perplexity levels. Notably, QEP-enhanced QuIP achieves state-of-the-art perplexity results among all tested layer-wise PTQ methods. Similar significant improvements are observed for RTN, GPTQ and AWQ at INT2g32, INT2g64, and INT2g128 quantization levels; see Appendix C.1 for details.
@@
We evaluate the zero-shot accuracy of quantized models on several tasks. Table 2 summarizes the average accuracy for the ArcE, PiQA, and SC datasets. Detailed results for each dataset are provided in Appendix C.2. Consistent with the perplexity results, QEP effectively improves existing layer-wise PTQ methods. Notably, the performance gains from QEP are especially pronounced with INT2 quantization. For Llama-2-70B, the QEP-enhanced QuIP at INT2 achieves performance comparable to RTN and GPTQ at INT3 quantization.
Table 3: Runtime comparison of the quantization process
@@
We examine the impact of computation time required to the correction term. Table 3 shows the processing time of each layer-wise PTQ. Since the quantization processing time for RTN is only a few seconds and thus negligible, the measured time for QEP+RTN is primarily due to computing the preprocessing of correction term. This result indicates that calculating the QEP correction term requires significantly less computation time compared to other existing layer-wise PTQ quantization process. Moreover, using the same calibration dataset for weight correction and quantization reduce preprocessing overhead by approximately one-half to one-third by reusing computational steps.
@@
As discussed in Section 5.3, our method adaptively controls propagation strength in Eq. (8) to mitigate overfitting to the calibration dataset. In this section, we empirically validate this approach. Table 4 compares the perplexity difference among QEP-enhanced RTN, GPTQ, and RTN when quantizing Llama-2-7B, evaluated on WikiText2 across various calibration datasets. Consistent with prior findings (Lin et al., 2024), GPTQ exhibits significant sensitivity to the calibration dataset: it outperforms RTN on C4 and WikiText2 but experiences notable performance degradation on PTB. In contrast, QEP-enhanced RTN consistently improves performance across all calibration datasets, demonstrating robustness to distributional shifts. This highlights the effectiveness of propagation control in preventing overfitting to the calibration datase.
@@ 7 Conclusion
Appendix A Additional Theoretical Results
This section provides proofs of the propositions stated in the main text and additional theoretical analyses. To avoid ambiguities, we fix the following notation throughout this section:
@@ Appendix A Additional Theoretical Results
Throughout the remainder of this section, we focus on first order expansion in the weight-quantization error. Specifically, we keep only linear terms in these errors and systematically discard all products that contain two of them:
${\bm{E}}_{i}{\bm{E}}_{j}=0,~{}~{}\forall i,j\in[L].$ (10)
This matches practice: with INT8 rounding $\|{\bm{E}}_{l}\|_{F}/\|{\bm{W}}_{l}\|_{F}=10^{-2}\sim 10^{-1}$, so quadratic terms are one order of magnitude smaller than any first–order contribution. Furthermore, The baseline PTQ and the QEP pipeline use the same quantiser configuration, hence they induce errors of the same order. We therefore write the same symbol ${\bm{E}}_{l}$ for the error matrix in either scheme; any difference is at most a few percent and does not affect first–order bounds.
A.1 Derivation of Proposition 5.1
@@
First, we rewrite the residual inside the Frobenius norm by using the following relationship: ${\bm{W}}_{l}{\bm{X}}_{l}={\bm{W}}{l}\widehat{{\bm{X}}}_{l}+{\bm{W}}_{l}{\bm{\delta}}_{l}$. Thus, the objective can be expressed as follows:
@@
To find the minimizer $\widehat{{\bm{W}}}_{l}$, we set the gradient of the expression with respect to $\widehat{{\bm{W}}}_{l}$, equal to zero. Using standard matrix calculus, we find that the calculus for a stationary point is
@@
A.2 Quantization Error Accumulation
This section demonstrates that under standard layer-wise PTQ, in which each layer is quantized independently without accounting for downstream effects, the activation difference at the output layer, defined as ${\bm{\delta}}_{L}\coloneqq{\bm{X}}_{L}-\widehat{{\bm{X}}}_{L}$, grows exponentially with depth, to first order in the quantization noise, under mild conditions.
Proposition A.1.
@@ A.2 Quantization Error Accumulation
${\bm{\delta}}_{l}=-\sum_{k=1}^{l-1}\left(\prod_{s=k+1}^{l-1}{\bm{J}}_{s}{\bm{W}}_{s}\right){\bm{J}}_{k}{\bm{E}}_{k}{\bm{X}}_{k}+\mathcal{O}({\bm{E}}_{1}^{2},{\bm{E}}_{1}{\bm{E}}_{2},\ldots,{\bm{E}}_{l}^{2}),$ (16)
@@
${\bm{X}}_{l}=\sigma_{l-1}({\bm{W}}_{l-1}{\bm{X}}_{l-1}),\quad\widehat{{\bm{X}}}_{l}=\sigma_{l-1}(\widehat{{\bm{W}}}_{l-1}\widehat{{\bm{X}}}_{l-1}).$ (17)
@@
${\bm{\delta}}_{l}$ $={\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$
$=\sigma_{l-1}({\bm{W}}_{l-1}{\bm{X}}_{l-1})-\sigma_{l-1}(\widehat{{\bm{W}}}_{l-1}\widehat{{\bm{X}}}_{l-1})$
$={\bm{J}}_{l-1}({\bm{W}}_{l-1}{\bm{X}}_{l-1}-\widehat{{\bm{W}}}_{l-1}\widehat{{\bm{X}}}_{l-1})+\mathcal{O}({\bm{E}}_{l-1}^{2},{\bm{\delta}}_{l-1}^{2})$
$={\bm{J}}_{l-1}\left[-{\bm{E}}_{l-1}{\bm{X}}_{l-1}-{\bm{W}}_{l-1}{\bm{\delta}}_{l-1}\right]+\mathcal{O}({\bm{E}}_{l-1}^{2},{\bm{\delta}}_{l-1}^{2})$
$=-{\bm{J}}_{l-1}{\bm{E}}_{l-1}{\bm{X}}_{l-1}-{\bm{J}}_{l-1}{\bm{W}}_{l-1}{\bm{\delta}}_{l-1}+\mathcal{O}({\bm{E}}_{l-1}^{2},{\bm{\delta}}_{l-1}^{2}).$
Expanding ${\bm{\delta}}_{l-1}$ explicitly, we obtain
${\bm{\delta}}_{l-1}=-{\bm{J}}_{l-2}{\bm{E}}_{l-2}{\bm{X}}_{l-2}-{\bm{J}}_{l-2}{\bm{W}}_{l-2}{\bm{\delta}}_{l-2}+\mathcal{O}({\bm{E}}_{l-2}^{2},{\bm{\delta}}_{l-2}^{2}).$ (18)
Substituting this into the previous expression explicitly yields
${\bm{\delta}}_{l}$ $=-{\bm{J}}_{l-1}{\bm{E}}_{l-1}{\bm{X}}_{l-1}$
$\quad-{\bm{J}}_{l-1}{\bm{W}}_{l-1}\left[-{\bm{J}}_{l-2}{\bm{E}}_{l-2}{\bm{X}}_{l-2}-{\bm{J}}_{l-2}{\bm{W}}_{l-2}{\bm{\delta}}_{l-2}\right]+\mathcal{O}({\bm{E}}_{l-1}^{2},{\bm{E}}_{l-2}^{2},{\bm{\delta}}_{l-2}^{2})$
$=-{\bm{J}}_{l-1}{\bm{E}}_{l-1}{\bm{X}}_{l-1}-{\bm{J}}_{l-1}{\bm{W}}_{l-1}{\bm{J}}_{l-2}{\bm{E}}_{l-2}{\bm{X}}_{l-2}$
$\quad-{\bm{J}}_{l-1}{\bm{W}}_{l-1}{\bm{J}}_{l-2}{\bm{W}}_{l-2}{\bm{\delta}}_{l-2}+\mathcal{O}({\bm{E}}_{l-1}^{2},{\bm{E}}_{l-2}^{2},{\bm{\delta}}_{l-2}^{2}).$
Repeating this explicit expansion recursively down to the first layer noting ${\bm{\delta}}_{1}={\bm{0}}$, we obtain the fully expanded form as follows:
${\bm{\delta}}_{l}=-\sum_{k=1}^{l-1}\left(\prod_{s=k+1}^{l-1}{\bm{J}}_{s}{\bm{W}}_{s}\right){\bm{J}}_{k}{\bm{E}}_{k}{\bm{X}}_{k}+\mathcal{O}({\bm{E}}_{1}^{2},{\bm{E}}_{1}{\bm{E}}_{2},\dots,{\bm{E}}_{l-1}^{2}),$ (19)

### v3

@@ 5.3 Controlling Propagation Strength
Although solving Eq. (7) effectively reduces the accumulation of quantization error, it can lead to overfitting. This issue is particularly pronounced when the calibration dataset is small and insufficiently representative of the target task, or when the model includes blocks with a large number of parameters such as the MLP blocks commonly found in transformer architectures, causing the correction to overfit the calibration dataset.
@@
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}}\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}^{2}+\lambda_{l}\|{\bm{W}}_{l}-\widehat{{\bm{W}}}_{l}\|_{F}^{2},~~\lambda_{l}\in{\mathbb{R}}_{+}.$ (11)
Specifically, as $\alpha_{l}$ increases from 0 to $1$, the corresponding parameter $\lambda_{l}$ decreases from $\infty$ to 0.
The derivation is provided in Appendix B.4. Additionally, the following proposition is established.
@@
Explicit conditions and comprehensive proofs of this proposition are provided in Appendix B.3. Consequently, the parameter $\alpha_{l}$ effectively controls overfitting, analogous to regularization techniques, and importantly provides a systematic way to balance overfitting and underfitting in layer-wise PTQ methods. Indeed, this parameter is crucial for preventing overfitting, especially in MLP blocks, which contain more parameters than other blocks.
Furthermore, in large-scale LLMs, the high-dimensional activations in MLP layers often result in computationally expensive correction terms. In these cases, selectively setting $\alpha_{l}=0$ for specific layers eliminates the computational cost of the correction term and acts as implicit regularization, potentially improving generalization. Therefore, appropriately setting $\alpha_{l}=0$ can reduce the correction time by approximately one-third and one-half. Developing adaptive strategies for layer-wise, data-aware, or resource-efficient tuning of $\alpha_{l}$ is a promising direction for future research. In the following, we refer to the overall approach, including the tunable mechanism controlled by $\{\alpha_{l}\}_{l=1}^{L}$, as Quantization Error Propagation (QEP).
@@ 6 Experiments
We use representative layer-wise PTQ methods based on linear quantization such as round-to-nearest (RTN) (frantar2022gptq; dettmers2023case), GPTQ (frantar2022gptq), AWQ (lin2024awq), and QuIP (chee2023quip). Although previous studies have explored extensions, such as non-linear and block-wise quantization, as discussed in Section 2, these techniques are orthogonal to the core improvement introduced by QEP. Therefore, to isolate and emphasize the impact of QEP, we focus on these representative layer-wise PTQ methods.
@@
This study focuses on weight-only quantization schemes, specifically per-channel and group-wise quantization, which have recently shown superior trade-offs between efficiency and accuracy (dettmers2023case; frantar2022gptq; lin2024awq). The main text evaluates per-channel quantization under INT4, INT3, and INT2 precision settings. Due to space constraints, detailed results for group-wise quantization are presented in Appendix D. For the propagation strength parameter $\alpha_{l}$, we adopt a representative default value of $\alpha_{l}=\nicefrac{{1}}{{2}}$ for all layers, except for the MLP layers in the Llama-2 70B model, for which we set $\alpha_{l}=0$. Tuning $\alpha_{l}$ can further improve performance but is beyond the scope of this study and is left for future work.
@@
Following previous studies, we evaluate the Hessian matrix using the same default calibration datasets used in their original implementations. Specifically, GPTQ and QuIP use the C4 dataset (frantar2022gptq) for calibration, while AWQ uses the Pile dataset (gao2020pile). Following frantar2022gptq, we evaluate the correction term in Eq. (6) using $128$ randomly sampled segments of $2048$ tokens each from the C4 dataset(raffel2020exploring), which consists of web-crawled text excerpts.
@@
Following lin2024awq; frantar2022gptq, we evaluate our method on recent popular LLMs, namely the Llama-2 and Llama-3 model families (touvron2023llama), with size ranging from 7B to 70B parameters, as well as Mistral-7 B (jiang2024identifying). These models demonstrate superior performance compared to other open-source LLMs (zhang2022opt; scao2022bloom) and have become widely adopted as foundational models for numerous derivative open-source models (taori2023stanfordalpaca; vicuna2023).
@@
Following established evaluation protocols from prior studies (dettmers2022gpt3; xiao2023smoothquant; frantar2022gptq; dettmers2023case; yao2022zeroquant), we evaluate the quantized LLMs using the perplexity (PPL) on WikiText2 (merity2016pointer), Penn Treebank (PTB) (marcus-etal-1994-penn), and C4 (raffel2020exploring), and zero-shot accuracy on benchmarks including ARC Easy (ArcE) (Boratko2018), PiQA (bisk2020piqa), and StoryCloze (SC) (mostafazadeh-etal-2016-corpus). Due to space limitations, detailed results for each dataset are provided in Appendix D. All experiments are conducted using a single NVIDIA V100 GPU.
@@ 6.1 Results
Table 1 summarizes PPL results of various quantized models evaluated on WikiText2, comparing several bit-widths and different layer-wise PTQ methods, both with and without QEP. Additional C4 and PTB dataset results are provided in Appendix D.1, demonstrating consistent trends in the following. Our results indicate that incorporating QEP significantly enhances the performance of layer-wise PTQ, substantially reducing perplexity across nearly all tested methods and quantization levels. In medium-bit scenarios such as INT4 and INT3, where AWQ already exhibits strong performance, applying QEP yields further improvements. At 2-bit quantization, existing layer-wise PTQ methods based on linear quantization typically suffer severe PPL degradation, rendering deployment infeasible. However, QEP effectively mitigates this issue, making INT2 quantization achievable with practical perplexity levels. Notably, QEP-enhanced QuIP achieves state-of-the-art perplexity results among all tested layer-wise PTQ methods. Similar significant improvements are observed for RTN, GPTQ, and AWQ at INT2g32, INT2g64, and INT2g128 quantization levels; see Appendix D.1 for details.
@@
We evaluate the zero-shot accuracy of quantized models on several tasks. Table 2 summarizes the average accuracy for the ArcE, PiQA, and SC datasets. Detailed results for each dataset are provided in Appendix D.2. Consistent with the perplexity results, QEP effectively improves existing layer-wise PTQ methods. Notably, the performance gains from QEP are especially pronounced with INT2 quantization. For Llama-2-70B, the QEP-enhanced QuIP at INT2 achieves performance comparable to RTN and GPTQ at INT3 quantization.
Table 3: Runtime comparison of the quantization process.
@@
We examine the impact of computation time required for the correction term. Table 3 shows the processing time of each layer-wise PTQ. Since the quantization processing time for RTN is only a few seconds and thus negligible, the measured time for QEP+RTN is primarily due to computing the preprocessing of the correction term. This result indicates that calculating the QEP correction term requires significantly less computation time than other existing layer-wise PTQ quantization processes. Moreover, using the same calibration dataset for weight correction and quantization reduces preprocessing overhead by approximately one-half to one-third by reusing computational steps.
@@
As discussed in Section 5.3, our method adaptively controls propagation strength in Eq. (10) to mitigate overfitting to the calibration dataset. In this section, we empirically validate this approach. Table 4 compares the perplexity difference among QEP-enhanced RTN, GPTQ, and RTN when quantizing Llama-2-7B, evaluated on WikiText2 across various calibration datasets. Consistent with prior findings (lin2024awq), GPTQ exhibits significant sensitivity to the calibration dataset: it outperforms RTN on C4 and WikiText2 but experiences notable performance degradation on PTB. In contrast, QEP-enhanced RTN consistently improves performance across all calibration datasets, demonstrating robustness to distributional shifts. This highlights the effectiveness of propagation control in preventing overfitting to the calibration dataset.
@@ 7 Conclusion
Limitations
QEP relies on a small calibration set, as in other layer-wise PTQ approaches, which makes performance sensitive to data quality; however, it overfits less than comparable methods such as GPTQ and AWQ. The method also introduces a per-layer propagation-strength parameter $\alpha_{l}$; Although this parameter is tunable, a fixed value, e.g., $\alpha_{l}=\nicefrac{{1}}{{2}}$, works well in most cases, and automatic learning of $\alpha_{l}$ is left for future work.
Appendix A Additional Related Work
Quantization error mitigation
Defensive Quantization (DQ) [lin2019defensive] mitigates error accumulation by introducing an orthogonality penalty to the weights, which reduces the correlation-driven amplification of quantization noise It also employs gradient-based quantization to optimize the positioning of quantization levels, further suppressing propagated error and enhancing robustness. In contrast, the QEP module serves as a plug-and-play component for general layer-wise quantization, explicitly modulating propagation strength under standard linear, gradient-free quantization methods, such as GPTQ, AWQ, and QuIP.
Relation to GPTAQ
GPTAQ [li2025gptaq] optimizes the same local objective but is closely related to GPTQ, and therefore does not easily generalize to other layer-wise PTQ methods such as AWQ and QuIP. In contrast, QEP adds the correction term in Eq. (6) directly to the pre-trained weights, enabling plug-and-play use with diverse PTQ algorithms and ensuring strong performance in low-bit regimes. Moreover, whereas GPTAQ offers no guarantee that its local optimization reduces global quantization error, our analysis provides such a guaranty. QEP also introduces a per-layer propagation-strength parameter, $\alpha_{l}$, which mitigates the overfitting often observed with GPTQ and GPTAQ.
Extreme low-bit layer-wise PTQ
In the ultra-low precision regime, the layer-wise PTQ pipeline described in the main text, such as GPTQ and AWQ with standard linear quantizers, often becomes inadequate. As a result, many methods adopt alternative formalisms that introduce additional degrees of freedom beyond naive rounding to avoid catastrophic quality degradation. Representative examples include SVID-based 1-bit parameterizations [xu2024onebit], token-adaptive mixtures of scaling factors [jo2024binarymos], and structured sparsity designed for extreme quantization [dong2025stbllm]. In parallel, binary-factor formats decompose each weight matrix into bit-packed sign factors with lightweight diagonal scaling, allowing inference to be largely driven by efficient 1-bit kernels [boza2025dbf, lee2025littlebit]. In these formats, optimization is more challenging: strictly binary variables and highly non-smooth objectives can make straight-through estimators brittle [long2021learning, ichikawa2025high, yin2019understanding], motivating the use of discrete optimization or robust relaxation-based solvers. Recent advances in controlled continuous relaxations and annealing-style objectives have emphasized promising techniques such as QQA and iSCO [ichikawa2025continuous, ichikawa2025optimization, sun2023revisiting, ichikawa2024controlling]. A natural direction for future work is to examine how these extreme low-bit approaches interact with our error-propagation perspective. For instance, we should investigate whether combining QEP with binary-factor or sparsity-based methods can further suppress cross-layer error growth while maintaining the strong INT2 performance observed here, yielding larger gains as precision approaches 1 bit.
Appendix B Additional Theoretical Results
This section presents proofs of the propositions stated in the main text along with additional theoretical analyzes. To avoid ambiguity, we will fix the following notation throughout this section:
@@ Appendix B Additional Theoretical Results
Throughout this section, we examine the first-order linear term in the weight perturbations $\{{\bm{E}}_{l}\}$. Concretely, we define each quantity as ${\bm{A}}={\bm{A}}^{(0)}+{\bm{A}}^{(1)}+{\bm{R}}$, where ${\bm{A}}^{(1)}$ collects all terms linear in $\{{\bm{E}}_{l}\}$ and the remainder satisfies $\|{\bm{R}}\|={\mathcal{O}}(\max_{k}\|{\bm{E}}_{k}\|_{2}^{2})$ as $\max_{k}\|{\bm{E}}_{k}\|_{2}\to 0$. This matches practice: with INT8 rounding $\|{\bm{E}}_{l}\|_{F}/\|{\bm{W}}_{l}\|_{F}=10^{-2}\sim 10^{-1}$, so quadratic terms are one order of magnitude smaller than any first–order contribution. Furthermore, The baseline PTQ and the QEP pipeline use the same quantiser configuration, hence they induce errors of the same order. We therefore write the same symbol ${\bm{E}}_{l}$ for the error matrix in either scheme; any difference is at most a few percent and does not affect first–order bounds.
B.1 Derivation of Proposition 5.1
@@
First, we rewrite the residual inside the Frobenius norm by using the following relationship: ${\bm{W}}_{l}{\bm{X}}_{l}={\bm{W}}_{l}\widehat{{\bm{X}}}_{l}+{\bm{W}}_{l}{\bm{\delta}}_{l}$. Thus, the objective can be expressed as follows:
@@
Since the objective is a strictly convex quadratic function of $\widehat{{\bm{W}}}_{l}$ when $\widehat{{\bm{H}}}_{l}$ is invertible, the stationary point is the unique minimizer. To find the minimizer $\widehat{{\bm{W}}}_{l}$, we set the gradient of the expression with respect to $\widehat{{\bm{W}}}_{l}$, equal to zero. Using standard matrix calculus, we find that the calculus for a stationary point is
@@
B.2 Quantization Error Accumulation
This section demonstrates that, under standard layer-wise PTQ, where each layer is quantized independently without considering downstream effects, the activation difference at the output layer, defined as ${\bm{\delta}}_{L}\coloneqq{\bm{X}}_{L}-\widehat{{\bm{X}}}_{L}$, grows exponentially with depth, to first order in the quantization noise, under mild conditions.
Proposition B.1.
@@ B.2 Quantization Error Accumulation
${\bm{\delta}}_{l}=-\sum_{k=1}^{l-1}\left(\prod_{s=k+1}^{l-1}{\bm{J}}_{s}{\bm{W}}_{s}\right){\bm{J}}_{k}{\bm{E}}_{k}{\bm{X}}_{k}+\mathcal{O}\left(\max_{k\leq l-1}\|{\bm{E}}_{k}\|_{F}^{2}\right),$ (17)
@@
${\bm{X}}_{l}=\sigma_{l-1}({\bm{W}}_{l-1}{\bm{X}}_{l-1}),~~\widehat{{\bm{X}}}_{l}=\sigma_{l-1}(\widehat{{\bm{W}}}_{l-1}\widehat{{\bm{X}}}_{l-1}).$ (18)
@@
${\bm{\delta}}_{l}$ $={\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$ (19)
$=\sigma_{l-1}({\bm{W}}_{l-1}{\bm{X}}_{l-1})-\sigma_{l-1}(\widehat{{\bm{W}}}_{l-1}\widehat{{\bm{X}}}_{l-1})$ (20)
$={\bm{J}}_{l-1}({\bm{W}}_{l-1}{\bm{X}}_{l-1}-\widehat{{\bm{W}}}_{l-1}\widehat{{\bm{X}}}_{l-1})+\mathcal{O}\left(\max\{{\bm{E}}_{l-1}^{2},{\bm{\delta}}_{l-1}^{2}\}\right)$ (21)
$={\bm{J}}_{l-1}\left[-{\bm{E}}_{l-1}{\bm{X}}_{l-1}+{\bm{W}}_{l-1}{\bm{\delta}}_{l-1}\right]+\mathcal{O}\left(\max\{{\bm{E}}_{l-1}^{2},{\bm{\delta}}_{l-1}^{2}\}\right)$ (22)
$=-{\bm{J}}_{l-1}{\bm{E}}_{l-1}{\bm{X}}_{l-1}+{\bm{J}}_{l-1}{\bm{W}}_{l-1}{\bm{\delta}}_{l-1}+\mathcal{O}\left(\max\{{\bm{E}}_{l-1}^{2},{\bm{\delta}}_{l-1}^{2}\}\right).$ (23)
By explicitly expanding ${\bm{\delta}}_{l-1}$, we obtain
${\bm{\delta}}_{l-1}=-{\bm{J}}_{l-2}{\bm{E}}_{l-2}{\bm{X}}_{l-2}+{\bm{J}}_{l-2}{\bm{W}}_{l-2}{\bm{\delta}}_{l-2}+\mathcal{O}\left(\max\{{\bm{E}}_{l-2}^{2},{\bm{\delta}}_{l-2}^{2}\}\right).$ (24)
Substituting this expression into the previous equation yields
${\bm{\delta}}_{l}$ $=-{\bm{J}}_{l-1}{\bm{E}}_{l-1}{\bm{X}}_{l-1}$ (25)
$~~+{\bm{J}}_{l-1}{\bm{W}}_{l-1}\left[-{\bm{J}}_{l-2}{\bm{E}}_{l-2}{\bm{X}}_{l-2}+{\bm{J}}_{l-2}{\bm{W}}_{l-2}{\bm{\delta}}_{l-2}\right]+\mathcal{O}\left(\max\{{\bm{E}}_{l-1}^{2},{\bm{E}}_{l-2}^{2},{\bm{\delta}}_{l-2}^{2}\}\right)$ (26)
$=-{\bm{J}}_{l-1}{\bm{E}}_{l-1}{\bm{X}}_{l-1}-{\bm{J}}_{l-1}{\bm{W}}_{l-1}{\bm{J}}_{l-2}{\bm{E}}_{l-2}{\bm{X}}_{l-2}$ (27)
$~~+{\bm{J}}_{l-1}{\bm{W}}_{l-1}{\bm{J}}_{l-2}{\bm{W}}_{l-2}{\bm{\delta}}_{l-2}+\mathcal{O}\left(\max\{{\bm{E}}_{l-1}^{2},{\bm{E}}_{l-2}^{2},{\bm{\delta}}_{l-2}^{2}\}\right).$ (28)
By recursively repeating this explicit expansion down to the first layer, we obtain the fully expanded form as follows, noting ${\bm{\delta}}_{1}={\bm{0}}$:
${\bm{\delta}}_{l}=-\sum_{k=1}^{l-1}\left(\prod_{s=k+1}^{l-1}{\bm{J}}_{s}{\bm{W}}_{s}\right){\bm{J}}_{k}{\bm{E}}_{k}{\bm{X}}_{k}+\mathcal{O}\left(\max_{k\leq l-1}\|{\bm{E}}_{k}\|^{2}\right),$ (29)